If the line y=√3x cuts the curve x^4 +ax^2 +bxy +cx +dy +6 =0 at A, B, C and D then, OA.OB.OC.OD (where O being origin) is what ???

Grade:11

1 Answers

Arun
6 years ago
x4+√3ax3+√3bx2+cx+√3dx+6=0,,x1,x2,x3,x4,are the four roots,OA=sqrt(x1^2+y1^2)=sqrt(x1^2+3x1^2)=2x1,OB=2x2, OC=2x3, OD=2x4, OA.OB.OC.OD=16X1*X2*X3*X4=16*6=96